OpenStudy (anonymous):

Suppose you roll two dice. Find the probability of not rolling a sum of 6

OpenStudy (anonymous):

@MAli13chineta

OpenStudy (anonymous):

@Mandre

OpenStudy (anonymous):

@WolfMaster

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

@AdmiralAngel25 @abb0t @amistre64

OpenStudy (anonymous):

can u help me wolf i ain understandin

OpenStudy (anonymous):

shes asking for help. @WolfMaster

OpenStudy (anonymous):

0-0

OpenStudy (anonymous):

it's not like im interested in a wolf

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Nice, so I show up 3rd when you press a? :D

OpenStudy (anonymous):

yea ;)

OpenStudy (anonymous):

well 4th

OpenStudy (anonymous):

i'm way down in the b's but any ways who is smart and can help

geerky42 (geerky42):

@MAli13chineta 's explanation doesn't help?

OpenStudy (anonymous):

i know especially if ur not Einstein

OpenStudy (amistre64):

determine how many ways there is to get a sum of 6

OpenStudy (anonymous):

5 ways

OpenStudy (anonymous):

i mean 4

OpenStudy (amistre64):

lets chk that 15 24 33 42 51 yep 5 ways and how many outcomes are there entotal?

OpenStudy (anonymous):

i mean 5 again my b

OpenStudy (anonymous):

yea thts wut i got

OpenStudy (amistre64):

do we agree that there are 36 outcomes total?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

the the prob of actually rolling a sum of 6 is 5/36 since all probs add up to 1, the rest of the probs are what does NOT add to 6 P(not6) = 1 - 5/36

OpenStudy (anonymous):

5 ways so 5/36 possibly

OpenStudy (anonymous):

what do u mean 1-5/36

OpenStudy (amistre64):

subtract what is 6 from 1, to determine what is NOT 6

OpenStudy (amistre64):

what is 1 - 5/36

OpenStudy (anonymous):

0.86

OpenStudy (amistre64):

P(6) + P(not6) = 1 therefore P(not6) = 1 - P(6)

OpenStudy (amistre64):

about .8611 yeah

OpenStudy (amistre64):

31/36 to be precise

OpenStudy (anonymous):

thx

OpenStudy (amistre64):

youre welcome